A balloon with a mass of 0.4 kg and volume of 0.5 m3 is released. What is the total force on the balloon if the density of the air is 1.2 kg / m3?
Answer #1

total force on the balloon = Fbyount - mg = V*(density of the air)*g - mg

= 0.5*1.2*9.8 - 0.4*9.8 = 1.96 N upward direction